# Quasinormal Modes and Stability Analysis of the JMN-1 Naked Singularity

### v1

@@ VI Discussions and Conclusions
To assess the stability of this naked singularity, we studied its behavior under linear axial perturbations such as scalar, electromagnetic, and gravitational. Then we have extracted the QNMs frequencies, which are the eigenmodes of the perturbations. By employing the methods of classical perturbation theory, we expanded the metric and retained only the linear terms to obtain the master equation. We impose appropriate boundary conditions as discussed in Sec.(III), namely an inner boundary condition at $r\to M_{0}R_{b}$ and an outer boundary condition at $r\to 100M$, which is comparable to spatial infinity Kokkotas:1999bd ; Berti:2009kk . While considering quantum gravity effects near singularity, we found that ingoing waves are absorbed at the inner boundary, and outgoing waves escape to infinity at the outer boundary.
@@
The potential of JMN-1 naked singularity is well-behaved and features a barrier structure that temporarily traps the incoming and outgoing waves. Also, the function $Q(x,\omega)$ has two turning points. Therefore, we applied the WKB method to solve the master equation and compute the QNM frequencies $\omega$. These frequencies are complex in nature: the real part corresponds to the oscillation frequency, while the imaginary part determines the damping rate. Over the years, the WKB method has been extensively developed to solve such differential equations in the context of gravitational perturbation theory. We found that our results are best captured by the third and sixth orders approximations, whereas deviations occur at higher order approximations. We observe that our WKB approach at these orders demonstrates strong convergence when we consider higher orders such as $7^{th}$ and $8^{th}$ (see Fig. 2). This clearly suggests that the WKB method is highly effective within a certain range of approximations. Moreover, we shoud note that in some cases much higher-order terms can lead to numerical instabilities or yield diminishing improvements in accuracy hatsuda2020 ; Zhao:2024 .
@@
It is worth noting that we did not compute the QNM frequencies for the fundamental mode $l=n=0$, as the WKB method is only valid when $l>n$ and does not yield precise results for the fundamental mode. Nevertheless, our results demonstrate that the imaginary part of the QNM frequencies remains negative, indicating that the JMN-1 naked singularity is stable under linear perturbations. Furthermore, our numerical simulations confirm the stability of the spacetime under scalar $(s=0)$, electromagnetic $(s=1)$, and gravitational $(s=2)$ perturbations. This is confirmed from tables 1,2, 3, and 4 where the imaginary part of the QNM frequency $\omega$ remains negative, indicating damping rather than exponential growth.
@@
Note that, we have also explored the parameter regime $34/5>M_{0}>2/$ 3 and $R_{b}\leq 3M$ of the JMN-1 naked singularity, as this is where a photon sphere is present and shadow formation occurs Saurabh:2023otl . This analysis is crucial for ensuring the stability of the JMN-1 naked singularity in the presence of a photon sphere and is also significant in the context of future EHT observations EventHorizonTelescope:2022xqj .
@@
In summary, we conclude that the JMN-1 naked singularity is stable under linear perturbations of scalar, electromagnetic, and gravitational nature. Future research could focus on investigating the implications of this stability in the context of observational constraints and gravitational wave signatures. We also plan to extend this stability analysis to check for the polar perturbations (even-parity) as well.

### v2

@@ VI Discussions and Conclusions
To assess the stability of this naked singularity, we studied its behavior under linear axial perturbations such as electromagnetic, and gravitational. Then we have extracted the QNMs frequencies, which are the eigenmodes of the perturbations. By employing the methods of classical perturbation theory, we expanded the metric and retained only the linear terms to obtain the master equation. We impose appropriate boundary conditions as discussed in Sec.(III), namely an inner boundary condition at $r\to R_{b}$ and an outer boundary condition at $r\to 100M$, which is comparable to spatial infinity Kokkotas:1999bd ; Berti:2009kk .
We would like to point out that the present analysis has been conducted for the exterior region of the JMN-1 spacetime. Specifically, we have focused on the vacuum solution extending from the matching radius up to spatial infinity as discussed above. This approach is consistent with existing literature on the computation of QNM frequencies of black holes, where the analysis is typically restricted to the exterior geometry, from the event horizon (e.g., $r=2M$ for Schwarzschild) to spatial infinity. The rationale behind this is that observational signatures, such as QNM frequencies, originate from perturbations in the exterior region, making this domain particularly relevant for gravitational wave detection. However, it is essential to note that investigating interior stability is equally important. We intend to address this in future work by incorporating quantum gravity-inspired models. In particular, we plan to explore effective geometries such as the Nariai-type universe or modifications arising from loop quantum gravity, which may provide a meaningful framework for analyzing the stability of the interior region and addressing the central singularity.
@@
The potential of JMN-1 naked singularity is well-behaved and features a barrier structure that temporarily traps the incoming and outgoing waves. In addition, the function $Q(x,\omega)$ has two turning points. Therefore, we applied the WKB method to solve the master equation and compute the QNM frequencies $\omega$. These frequencies are complex in nature: the real part corresponds to the oscillation frequency, while the imaginary part determines the damping rate. Over the years, the WKB method has been extensively developed to solve such differential equations in the context of gravitational perturbation theory. We found that our results are best captured by orders between 3rd and 6th, whereas deviations occur at higher order approximations. We observe that our WKB approach at these orders demonstrates strong convergence when we consider higher orders such as $7^{th}$ and $8^{th}$ (see Fig. 2). This clearly suggests that the WKB method is highly effective within a certain range of approximations. Moreover, we shoud note that in some cases much higher-order terms can lead to numerical instabilities or yield diminishing improvements in accuracy hatsuda2020 ; Zhao:2024 .
@@
It is worth noting that we did not compute the QNM frequencies for the fundamental mode $l=n=0$, as the WKB method is only valid when $l>n$ and does not yield precise results for the fundamental mode. Nevertheless, our results demonstrate that the imaginary part of the QNM frequencies remains negative, indicating that the JMN-1 naked singularity is stable under linear perturbations. Furthermore, our numerical simulations confirm the stability of the spacetime under electromagnetic $(s=1)$, and gravitational $(s=2)$ perturbations. This is confirmed from tables 1,2, 3, and 4 where the imaginary part of the QNM frequency $\omega$ remains negative, indicating damping rather than exponential growth.
@@
Note that we have explored the parameter regime $4/5>M_{0}>2/3$ and $R_{b}\leq 3M$ of the JMN-1 naked singularity, as this is where a photon sphere is present and shadow formation occurs Saurabh:2023otl . This analysis is crucial for ensuring the stability of the JMN-1 naked singularity in the presence of a photon sphere and is also significant in the context of future EHT observations EventHorizonTelescope:2022xqj .
@@
In summary, we conclude that the JMN-1 naked singularity is stable under axial linear perturbations. Future research could focus on investigating the implications of this stability in the context of observational constraints and gravitational wave signatures. It is important to emphasize that our present analysis has focused exclusively on the axial (odd-parity) sector, following the standard approach in perturbation theory where axial stability is often treated as the first step Regge:1957td ; Moncrief:1974ng ; Noui:2023ksf . A comprehensive stability assessment also requires the study of polar (even-parity) perturbations, which we plan to address in forthcoming work. Together, these analyses will provide a complete picture of the stability properties of the spacetime.